Find the solution of the initial value problem y – 2y + 5y = g(t), y(0) = 0, y(0) = 0, where g(t) is a continuous, otherwi
Answer #1

Y- 2yt sy= g(t), y(o)=0, y(o)=0. ㅋ 0 ing laplace Tar [s²yes)- sy col-yo)) - 2 (sycs)-ycod] + 5ycs) = G(s) putting given in